# DistillNeRF: Perceiving 3D Scenes from Single-Glance Images by Distilling Neural Fields and Foundation Model Features

### v1

@@ Abstract
We propose DistillNeRF, a self-supervised learning framework addressing the challenge of understanding 3D environments from limited 2D observations in autonomous driving. Our method is a generalizable feedforward model that predicts a rich neural scene representation from sparse, single-frame multi-view camera inputs, and is trained self-supervised with differentiable rendering to reconstruct RGB, depth, or feature images. Our first insight is to exploit per-scene optimized Neural Radiance Fields (NeRFs) by generating dense depth and virtual camera targets for training, thereby helping our model to learn 3D geometry from sparse non-overlapping image inputs. Second, to learn a semantically rich 3D representation, we propose distilling features from pre-trained 2D foundation models, such as CLIP or DINOv2, thereby enabling various downstream tasks without the need for costly 3D human annotations. To leverage these two insights, we introduce a novel model architecture with a two-stage lift-splat-shoot encoder and a parameterized sparse hierarchical voxel representation. Experimental results on the NuScenes dataset demonstrate that DistillNeRF significantly outperforms existing comparable self-supervised methods for scene reconstruction, novel view synthesis, and depth estimation; and it allows for competitive zero-shot 3D semantic occupancy prediction, as well as open-world scene understanding through distilled foundation model features. Demos and code will be available at https://distillnerf.github.io/.
@@ 1 Introduction
Neural scene representations, such as NeRFs [7, 8] and 3D Gaussian Splatting (3DGS) [9], have recently emerged as a compelling paradigm for learning 3D representations from 2D signals in a self-supervised manner. While these methods demonstrated strong capabilities in view synthesis for indoor scenes [10], and more recently also for challenging dynamic outdoor scenes [11, 12, 13, 14], they require training a separate representation for each new scene, leading to extensive computation needs, typically in the order of hours. Additionally, most works focus on view synthesis only [11, 12, 13], and do not fully leverage the profound capability of neural scene representations to lift various sources of 2D information to 3D, such as features from 2D visual foundation models [15, 16].
@@
To this end, we propose DistillNeRF (Fig. 1), a conceptually simple framework for learning generalizable neural scene representations for driving scenes, by distilling 1) offline optimized NeRFs for enhanced geometry and appearance, and 2) pre-trained 2D foundation models, DINOv2 [17, 16] and CLIP [15], for enriched semantics. Generalizable scene representations with feed-forward models is an active area of research [18, 19, 20]. The autonomous driving domain remains particularly challenging due to sparse camera views with little overlap. Indeed, prior generalizable NeRFs for driving are shown to be struggling with view synthesis [21, 22, 23]. In contrast, we show that by distilling per-scene optimized NeRFs and visual foundation models, DistillNeRF allows predicting a 3D feature volume with strong geometry, appearance, and semantics, from sparse single-timestep images. The representation is capable of rendering tasks without per-scene optimization (e.g. scene reconstruction, novel view synthesis, foundation model feature prediction), and support downstream tasks, such as open-vocabulary text query and zero-shot 3D semantic occupancy prediction.
@@
DistillNeRF comprises two stages: offline per-scene NeRF training, and distillation into a generalizable model. The first stage trains a NeRF for each scene individually from each driving log, exploiting all available multi-view, multi-timestep information. Specifically, we use EmerNerf [14], a recent NeRF approach with decomposed static and dynamic fields. The second stage trains a generalizable encoder to directly lift multi-camera 2D images captured at a single timestep to a 3D continuous feature field, from which we render images, and supervise with dense depth and novel-view RGB targets generated from the per-scene optimized NeRFs, and foundation model features. Specifically, we propose a novel model architecture with 1) a two-stage Lift-Splat-Shoot encoder [24] to lift 2D observations into 3D; 2) a sparse hierarchical 3D voxel for efficient runtime and memory, parameterized to account for unbounded driving scenes; 3) feature image generation via differentiable volumetric rendering, decoded into appearance, depth, and optionally, foundation model features.
@@
Extensive experiments on the NuScenes dataset [25] demonstrate that DistillNeRF allows for high-quality scene reconstruction and novel view synthesis in previously unseen environments, on par with test-time per-scene optimization approaches, and significantly outperforming previous generalizable approaches. We also show strong results for zero-shot 3D semantic occupancy prediction, and promising quantitative results for open-vocabulary scene understanding.
@@
We summarize our main contributions as follows:
  • We introduce a conceptually simple but highly capable approach, named DistillNeRF, which enables a range of view synthesis and 3D-related scene understanding tasks by distilling from offline per-scene optimized NeRFs and 2D foundation models.
  • We propose a novel generalizable NeRF model architecture with sparse hierarchical voxel representation, which lifts and fuses multi-view 2D image features into a 3D parameterized neural field, and uses volumetric rendering for self-supervised training.
  • We conduct extensive experiments to demonstrate DistillNeRF’s strong performance on rendering and downstream tasks.
@@ 2 Related Work
Neural scene representations, like NeRFs [7, 26] and 3DGS [9], have brought unprecedented success in learning powerful representations of 3D scenes, and have also been successfully applied to challenging driving scenes populated with dynamic objects [27, 28, 12, 29, 30, 31, 32, 33]. However, these methods typically require expensive training for each scene, typically in the order of hours.
@@
Generalizable Neural Radiance Fields, such as [18, 34, 35, 36, 37], adapt the capabilities of conventional NeRFs for 3D scene reconstruction and novel view synthesis into a generalizable feedforward model. They replace the costly per-scene optimization with a single feedforward pass through the models. Recent works have extended such approaches to challenging driving scenes and demonstrate the potential in down-stream tasks [22, 5, 23, 38, 39]. However, due to the challenging sparse-view limited-overlapping camera settings on vehicles, these methods usually fail to show strong rendering performance. To the best of our knowledge, we are the very first work to achieve strong scene reconstruction and reasonable novel-view synthesis in driving scenes.
@@
Recent advancements extend NeRFs beyond novel view synthesis by integrating 2D features from foundation models into 3D space, equipping neural fields with semantic understanding [40]. Recent approaches also demonstrate similar capabilities in outdoor driving scenes [14] by distilling DINO features into the scene representation. However, these approaches suffer from prolonged optimization times when combined with feature distillation, and thus are impractical for online autonomous driving due to their costly per-scene optimization. Closely related to our work, FeatureNeRF [20] is a generalizable method that distills DINOv2 [17] features into 3D for keypoint transfer, but only investigates simple synthetic datasets like ShapeNet [41]. In contrast, we address the more challenging setting of autonomous driving. Our method, DistillNeRF, infers 3D feature fields in a single forward pass, making real-time application possible.
@@
Model distillation is a well-established idea [42]. NeRFs have also been distilled into, e.g., Generative Adversarial Networks in [43], and feed-forward models for temporal object shape prediction in [44]. However, prior work mainly focuses on static, object-centric, or indoor scenes. To the best of our knowledge, we are the first to propose distilling a per-scene optimized NeRFs with static-dynamic decomposition into a generalizable model for driving scenes.
@@ 3 Method
DistillNeRF predicts a generalizable scene representation in the form of sparse hierarchical voxels from single-timestep multi-view RGB image inputs, and is trained through volumetric rendering to output RGB, depth, and feature images.
The method is depicted in Fig. 1, the detailed architecture in Fig. 2, and key capabilities in Fig. 3. Inputs are $N$ posed RGB camera images $\{I_{i}\}^{N}_{i=1}$. We use a 2D backbone to extract $N$ feature images $\{X_{i}\}_{i=1}^{N}$. We then lift the 2D features to a 3D voxel-based neural field $\mathcal{V}\in\mathbb{R}^{H\times W\times D\times C}$ using the corresponding camera matrix, and apply sparse quantization and convolution to fuse features from multiple views. To account for unbounded scenes we use a parameterized neural field with fixed-scale inner voxels, and varying-scale outer voxels contracting the infinite range. Volumetric rendering is performed to supervise the reconstruction of the scene. For better guidance on scene geometry, we “distill” knowledge from offline optimized NeRFs, using rendered dense depth images from original camera views and virtual camera views. Foundation model features, from CLIP or DINOv2, are set as additional reconstruction objectives and thus are also “distilled” into our model to enrich scene semantics. We introduce design principals in the following section, and refer to Appendix A.1 for details.
@@ 3.1 Sparse Hierarchical Voxel Model
For each of the $N$ camera image inputs, we follow a similar procedure as Lift-Splat-Shoot (LSS) [24] to lift the 2D image features to the 3D neural field. Unlike typical LSS and variants [24, 39, 45] that predict depth in one shot, we propose a two-stage, coarse-to-fine strategy with two jointly trained predictors to capture more nuanced depth. Following prior works, the first stage predicts categorical depth and aggregates them into a single prediction with ray marching. The second stage then predicts a distribution over a fine-grained set of categorical depth values, chosen around the coarse depth prediction.
Specifically, in the first stage, we feed each image to a 2D backbone to generate a depth feature map of size $H\times W\times D$. Inspired by the volume rendering equation [7], the depth feature map is regarded as a discrete frustum where $D$ denotes the number of pre-defined categorical depths. Each entry in the frustum is a density value. That is, the $d$’th channel of the frustum at pixel $(h,w)$ represents the density value $\sigma_{h,w,d}$ of the frustum entry at $(h,w,d)$. The occupancy weight of entry $(h,w,d)$ is then
@@
In the second stage, we dynamically generate a fine-grained set of $D^{\prime}$ depth candidates centered around the coarse depth prediction. We then combine learned embeddings for each candidate with the depth features from the first stage, and feed them to another network to generate the density of each depth candidate. The occupancy weights $\mathbb{O}^{\prime}$ of the fine-grained depth candidates are predicted similarly by Eq 1, which can also be regarded as probabilities of each depth candidate.
With the candidate depths associated with probabilities, we then lift 2D image features to 3D. Specifically, we use the feature pyramid network (FPN) [46] to get 2D image features $\phi$, and assign the 2D image features to the 3D frustum. That is, for pixel $(h,w)$, its image feature $\phi_{h,w}$ is distributed to each depth candidates $t^{\prime}_{d}$ by $[\mathbb{O}^{\prime}_{h,w,d}\phi_{h,w},\sigma^{\prime}_{h,w,d}]$, where we scale the pixel image feature $\phi_{h,w}$ with occupancy $\mathbb{O}^{\prime}_{h,w,d}$ and concatenate it with density $\sigma^{\prime}_{h,w,d}$.
Figure 3: Given single-frame multi-view cameras as input and without test-time per-scene optimization, DistillNeRF can reconstruct RGB images (row 2), estimate depth (row 3), render foundation model features (rows 4, 5) which enables open-vocabulary text queries (rows 6, 7, 8), and predict binary and semantic occupancy in zero shot (rows 9, 10).
@@
After constructing the frustum for each view, we transform the frustums to the world coordinates using the camera extrinsics, and fuse them into a shared 3D voxel-based neural field $\mathcal{V}$, where each voxel represents a region in the world coordinates and carries both densities and features. When lifted frustum entries from different views lie in the same voxel, we fuse them with average pooling.
@@
Unlike previous works [39, 47, 21] using dense voxels, which uniformly quantizes the neural field and potentially wastes computation and memory on large empty regions, we apply sparse quantization on the neural field. Specifically, we follow the octree representation [48] to recursively divide the neural field into specified levels, according to the 3D positions of the lifted 2D features. While an octree with many levels and thus smaller voxel sizes can capture more accurate 3D positions of lifted features, overly fine-grained octrees can lead to difficulty in querying features during rendering (e.g. missing far-away features due to large gaps between sampled rays.) To this end, we generate two octrees with different quantization granularities, one fine octree with more quantization levels capturing details of the lifted features, and one coarse octree to represent general information of a larger range. Sparse convolutions [49] are then applied to both octrees to encode the relationships and interactions among voxels. The features in the fine octree are also downsampled and concatenated with the coarse octree to enhance details.
@@
Unlike prior works that consider a neural field covering a fixed range [39, 21, 22], our work aims at accounting for the unbounded-scene settings in the driving scenes by proposing a parameterized neural field. We want to keep the inner-range voxels at the real scale and high resolution due to their importance to various tasks (e.g., occupancy prediction in 50 meters’ range), while contracting the scene up to infinite distance in the outer range of the voxels at a lower resolution, so we can render with low memory and computation cost (e.g. sky, far-away buildings). Inspired by [8, 50], we propose a transform function that maps a 3D point in the world coordinates $p=(x,y,z)$ to the coordinates in the parameterized neural field:
@@
The transformed coordinates $f(p)$ will always be within $[0,1]$, where $p_{inner}$ sets the range of the inner voxel (region of interest) and varies in $x,y,z$ directions, and $\alpha\in[0,1]$ denotes the proportion of the inner range in the parameterized neural field. Consistent parameterizations are enforced for both the single-view lifting process (on the depth space) and the multi-view fusion process (on the 3D coordinate space).
@@
Finally, we use differentiable volumetric rendering to project the 3D neural field onto 2D feature maps and render images. Specifically, for each pixel of each camera, we shoot a ray originating from the camera to the neural field according to the camera poses, and sample points along the ray. For each sample point, we query both fine and coarse octree to get the density and features of the corresponding voxel that the sample point lies in. Further, to capture both high-level information and fine-grained details, the features from both octrees are concatenated as the final feature. Regarding the density, while the fine octree captures more accurate 3D positioning, since the fine octree voxel only covers a small region, the sample points could be easily within empty voxels and thus query no information, especially for faraway regions. To this end, for each sample point, we first query the fine octree to get the fine density. If the fine density is empty, we query the coarse octree to complement the density. Regarding tne sampling strategy, we follow [8] to sample points for each ray with two phases: first we sample a set of points uniformly, then we sample another set of points with importance sampling given densities for the first set of points, so to enhance surface details in the scene. With the densities and features of the sampled points we do volumetric rendering using Eq 2 to get the 2D feature map for each camera. The rendered 2D feature maps are then fed into a CNN decoder to enhance high-frequency details, and upsample the final RGB image without increasing the rendering resolution/cost. Note that from the volume rendering process, we can also get the expected depth for each pixel [27].

### v2

@@ Abstract
We propose DistillNeRF, a self-supervised learning framework addressing the challenge of understanding 3D environments from limited 2D observations in outdoor autonomous driving scenes. Our method is a generalizable feedforward model that predicts a rich neural scene representation from sparse, single-frame multi-view camera inputs with limited view overlap, and is trained self-supervised with differentiable rendering to reconstruct RGB, depth, or feature images. Our first insight is to exploit per-scene optimized Neural Radiance Fields (NeRFs) by generating dense depth and virtual camera targets from them, which helps our model to learn enhanced 3D geometry from sparse non-overlapping image inputs. Second, to learn a semantically rich 3D representation, we propose distilling features from pre-trained 2D foundation models, such as CLIP or DINOv2, thereby enabling various downstream tasks without the need for costly 3D human annotations. To leverage these two insights, we introduce a novel model architecture with a two-stage lift-splat-shoot encoder and a parameterized sparse hierarchical voxel representation. Experimental results on the NuScenes and Waymo NOTR datasets demonstrate that DistillNeRF significantly outperforms existing comparable state-of-the-art self-supervised methods for scene reconstruction, novel view synthesis, and depth estimation; and it allows for competitive zero-shot 3D semantic occupancy prediction, as well as open-world scene understanding through distilled foundation model features. Demos and code will be available at https://distillnerf.github.io/.
@@ 1 Introduction
Neural scene representations, such as NeRFs [7, 8] and 3D Gaussian Splatting (3DGS) [9], have recently emerged as a compelling paradigm for learning 3D representations from 2D signals in a self-supervised manner. While these methods demonstrated strong capabilities in view synthesis for indoor scenes [10], and more recently also for challenging dynamic outdoor scenes [11, 12, 13, 14], they require delicately training a new representation for each new scene, leading to extensive computation budget and time needs, typically in the order of hours or minutes. This falls short of the real-time computational requirements for autonomous driving, which typically demand a processing speed of 2-10 Hz. Additionally, most works focus on view synthesis only [11, 12, 13], resulting in learned 3D representations that lack semantics, sidestep downstream tasks, and do not fully exploit the potential of neural scene representations to transform various sources of 2D information into 3D, such as features from 2D visual foundation models [15, 16].
@@
To this end, we propose DistillNeRF (Fig. 1), a conceptually simple framework for learning generalizable neural scene representations in driving scenes, by distilling 1) offline optimized NeRFs for enhanced geometry and appearance, and 2) pre-trained 2D foundation models, DINOv2 [17, 16] and CLIP [15], for enriched semantics. Generalizable scene representations with feed-forward models is an active area of research [18, 19, 20], and the autonomous driving domain remains particularly challenging due to sparse camera views with little overlap: prior generalizable NeRFs for driving are shown to be struggling with view synthesis [21, 22, 23]. In contrast, we show that by distilling per-scene optimized NeRFs and visual foundation models, DistillNeRF allows predicting a 3D feature volume with strong geometry, appearance, and semantics, from sparse single-timestep images. The representation is capable of rendering tasks without per-scene optimization (e.g. scene reconstruction, novel view synthesis, foundation model feature prediction), and support downstream tasks, such as open-vocabulary text query and zero-shot 3D semantic occupancy prediction.
@@
DistillNeRF comprises two stages: offline per-scene NeRF training, and distillation into a generalizable model. The first stage trains a NeRF for each scene individually from each scene’s driving log, exploiting all available multi-view, multi-timestep information. Specifically, we use EmerNerf [14], a recent NeRF approach with decomposed static and dynamic fields. The second stage trains a generalizable encoder to directly lift multi-camera 2D images captured at a single timestep to a 3D continuous feature field, from which we render images, and supervise with dense depth and novel-view RGB targets generated from the per-scene optimized NeRFs, and foundation model features. Specifically, we propose a novel model architecture with 1) a two-stage Lift-Splat-Shoot encoder [24] to lift 2D observations into 3D; 2) a sparse hierarchical 3D voxel for efficient runtime and memory, parameterized to account for unbounded driving scenes; 3) feature image generation via differentiable volumetric rendering, decoded into appearance, and optionally, foundation model features.
@@
Extensive experiments on the NuScenes [25] and Waymo NOTR [26, 14] dataset demonstrate that DistillNeRF allows for high-quality scene reconstruction and novel view synthesis in previously unseen environments without per-scene training, on par with test-time per-scene optimization approaches, and significantly outperforming previous generalizable approaches. We also show strong results for zero-shot 3D semantic occupancy prediction, and promising quantitative results for open-vocabulary scene understanding.
@@ 2 Related Work
Neural scene representations, like NeRFs [7, 27] and 3DGS [9, 28], have brought unprecedented success in learning powerful representations of 3D scenes, and have also been successfully applied to challenging driving scenes populated with dynamic objects [29, 30, 12, 31, 32, 33, 34, 35]. However, these methods typically require expensive training for each scene, typically in the order of hours or minutes.
@@
Generalizable Neural Radiance Fields, such as [18, 36, 37, 38, 39], adapt the capabilities of conventional NeRFs for 3D scene reconstruction and novel view synthesis into a generalizable feedforward model. They replace the costly per-scene optimization with a single feedforward pass through the models. Recent works have extended such approaches to challenging driving scenes and demonstrated the potential in down-stream tasks [22, 5, 23, 40, 41]. However, due to the challenging sparse-view limited-overlapping camera settings on vehicles, these methods usually fail to show strong rendering performance. To the best of our knowledge, we are the very first work to achieve strong scene reconstruction and reasonable novel-view synthesis on par with offline per-scene optimized NeRFs in driving scenes.
@@
Recent advancements extend NeRFs beyond novel view synthesis by integrating 2D features from foundation models into 3D space, equipping neural fields with semantic understanding [42]. Recent approaches also demonstrate similar capabilities in outdoor driving scenes [14] by distilling DINO features into the scene representation. However, these approaches suffer from prolonged optimization times when combined with feature distillation, and thus are impractical for online autonomous driving due to their costly per-scene optimization. Closely related to our work, FeatureNeRF [20] is a generalizable method that distills DINOv2 [17] features into 3D for keypoint transfer, but only investigates simple indoor object-level synthetic datasets like ShapeNet [43], where a large number of overlapping camera images are placed around the object to scan it from various angles. In contrast, we address the more challenging outdoor scene-level settings of autonomous driving, with sparse limited-overlapping camera image inputs. Our method, DistillNeRF, infers 3D feature fields in a single forward pass, making real-time application possible.
@@
Model distillation is a well-established idea [44]. NeRFs have also been distilled into, e.g., Generative Adversarial Networks in [45], and feed-forward models for temporal object shape prediction in [46]. However, prior work mainly focuses on static, object-centric, or indoor scenes. To the best of our knowledge, we are the first to propose distilling a per-scene optimized NeRFs with static-dynamic decomposition into a generalizable model for outdoor driving scenes.
@@ 3 Method
DistillNeRF predicts a generalizable scene representation in the form of sparse hierarchical voxels from single-timestep multi-view RGB image inputs, and is self-supervised by natural sensor streams, through volumetric rendering to output RGB, depth, and feature images.
The method is depicted in Fig. 1, the detailed architecture in Fig. 2, and key capabilities in Fig. 3 and Fig. 4. Inputs are $N$ posed RGB camera images $\{I_{i}\}^{N}_{i=1}$. We use a 2D backbone to extract $N$ feature images $\{X_{i}\}_{i=1}^{N}$. We then lift the 2D features to a 3D voxel-based neural field $\mathcal{V}\in\mathbb{R}^{H\times W\times D\times C}$, and apply sparse quantization and convolution to fuse features from multiple views. To account for unbounded scenes we use a parameterized neural field with fixed-scale inner voxels, and varying-scale outer voxels contracting the infinite range. Volumetric rendering is performed to supervise the reconstruction of the scene. For better guidance on scene geometry, we “distill” knowledge from offline optimized NeRFs, using rendered dense depth images from original camera views and virtual camera views. Foundation model features, from CLIP or DINOv2, are set as additional reconstruction objectives and thus are also “distilled” into our model to enrich scene semantics. We introduce design principals in the following section, and refer to Appendix A.7 for implementation details.
@@ 3.1 Sparse Hierarchical Voxel Model
For each of the $N$ camera image inputs, we follow a similar procedure as Lift-Splat-Shoot (LSS) [24] to lift the 2D image features to the 3D neural field. Unlike typical LSS and variants [24, 41, 47] that predict depth in one shot, we propose a two-stage, coarse-to-fine strategy with two jointly trained predictors to capture more nuanced depth. The first stage, following prior LSS works, predicts categorical depth and aggregates them into a single prediction with ray marching. The second stage then predicts a distribution over a fine-grained set of categorical depth values, which are centered around the first-stage predicted coarse depth.
Specifically, in the first stage, we feed each image to a 2D backbone to generate a depth feature map of size $H\times W\times D$. The depth feature map is regarded as a discrete frustum where $D$ denotes the number of pre-defined categorical depths. Inspired by the volume rendering equation [7], each entry in the frustum is a density value. That is, the $d$’th channel of the frustum at pixel $(h,w)$ represents the density value $\sigma_{h,w,d}$ of the frustum entry at $(h,w,d)$. The occupancy weight of entry $(h,w,d)$ is then
@@
In the second stage, centered around the initial coarse depth prediction, we dynamically sample a set of $D^{\prime}$ fine-grained depth candidates. This involves uniform sampling, with the sampling range adaptively adjusted based on the coarse depth estimate. We then embed these fine-grained depth candidates, and combine their embeddings with the depth features from the first stage, and feed them to another network to generate the density of each fine-grained depth candidate. The occupancy weights $\mathbb{O}^{\prime}$ of the fine-grained depth candidates are predicted similarly by Eq 1, which can also be regarded as probabilities of each fine-grained depth candidate.
With the candidate depths associated with probabilities, we then lift 2D image features to 3D. Specifically, we use a 2D image backbone to get 2D image features $\phi$, and assign the 2D image features to the 3D frustum according to each pixel’s depth. That is, for pixel $(h,w)$, its image feature $\phi_{h,w}$ is distributed to each fine-grained depth candidates $t^{\prime}_{d}$ by $[\mathbb{O}^{\prime}_{h,w,d}\phi_{h,w},\sigma^{\prime}_{h,w,d}]$, where we scale the pixel image feature $\phi_{h,w}$ with occupancy $\mathbb{O}^{\prime}_{h,w,d}$ and concatenate it with density $\sigma^{\prime}_{h,w,d}$.
Figure 3: DistillNeRF Capabilities - Given single-frame multi-view cameras as input and without test-time per-scene optimization, DistillNeRF can reconstruct RGB images (row 2), estimate depth (row 3), render foundation model features (rows 4, 5) which enables open-vocabulary text queries (rows 6, 7, 8), and predict binary and semantic occupancy in zero shot (rows 9, 10).
@@
After constructing the frustum for each view, we transform the frustums to the world coordinates using the camera poses, and fuse them into a shared 3D voxel-based neural field $\mathcal{V}$, where each voxel represents a region in the world coordinates and carries both densities and features. When lifted frustum entries from different views lie in the same voxel, we fuse them with average pooling.
@@
Unlike previous works [41, 48, 21] using dense voxels, which uniformly quantizes the neural field and potentially wastes computation and memory on large empty regions, we apply sparse quantization on the neural field. Specifically, we follow the octree representation [49] to recursively divide the neural field into specified levels, according to the 3D positions of the lifted 2D features. While an octree with many levels and thus smaller voxel sizes can capture more accurate 3D positions of lifted features, overly fine-grained octrees can lead to difficulty in querying features during rendering (e.g. missing far-away features due to large gaps between sampled rays). To this end, we generate two octrees with different quantization granularities, one fine octree with more quantization levels capturing details of the lifted features, and one coarse octree to represent general information of a larger range. Sparse convolutions [50] are then applied to both octrees to encode the relationships and interactions among voxels, during which the features in the fine octree are also downsampled and concatenated with the coarse octree to enhance details.
@@
Unlike prior works that consider a neural field covering a fixed range [41, 21, 22], our work aims at accounting for the unbounded-scene settings in the driving scenes by proposing a parameterized neural field. We want to keep the inner-range voxels at the real scale and high resolution due to their importance to various downstream tasks (e.g., occupancy prediction in 50 meters’ range), while contracting the scene up to infinite distance in the outer range of the voxels at a lower resolution, so we can render with low memory and computation cost (e.g. sky, far-away buildings). Inspired by [8, 51], we propose a transform function that maps a 3D point in the world coordinates $p=(x,y,z)$ to the coordinates in the parameterized neural field:
@@
The transformed coordinates $f(p)$ will always be within $[0,1]$, where $p_{inner}$ sets the range of the inner voxel (region of interest) and varies in $x,y,z$ directions, and $\alpha\in[0,1]$ denotes the contraction ratio, namely the proportion of the inner range in the parameterized neural field. Consistent parameterizations are enforced for both the single-view lifting process (on the depth space) and the multi-view fusion process (on the 3D coordinate space).
Figure 4: DistillNeRF Generalizability - Trained on the nuScenes dataset, our model demonstrates strong zero-shot transfer performance on the unseen Waymo NOTR dataset, achieving decent reconstruction quality (row 2). This quality can be further enhanced by applying simple color alterations to account for camera-specific coloring discrepancies (row 3). After fine-tuning (row 4), our model surpasses the offline per-scene optimized EmerNeRF, achieving higher PSNR (29.84 vs. 28.87) and SSIM (0.911 vs. 0.814). See Tab 3 for quantitative results.
@@
Finally, we use differentiable volumetric rendering to project the 3D neural field onto 2D feature maps and render images. Specifically, for each pixel of each camera, we shoot a ray originating from the camera to the neural field according to the camera poses, and sample points along the ray. Feature Querying: For each sample point, we query both fine and coarse octree to get the density and features of the corresponding voxel that the sample point lies in. Further, to capture both high-level information and fine-grained details, the features from both octrees are concatenated as the final feature. Density Querying: Regarding the density, while the fine octree captures more accurate 3D positioning, since the fine octree voxel only covers a small region, the sample points could be easily within empty voxels and thus query no information, especially for faraway regions. To this end, for each sample point, we first query the fine octree to get the fine density. If the fine density is empty, we query the coarse octree to complement the density. Two-Stage Sampling: Regarding the sampling strategy, we follow [8] to sample points for each ray with two phases: first we sample a set of points uniformly, then we sample another set of points with importance sampling given densities for the first set of points, so to enhance surface details in the scene. With the densities and features of the sampled points we do volumetric rendering using Eq 2 to get the 2D feature map for each camera. Decoding: The rendered 2D feature maps are then fed into a CNN decoder to enhance high-frequency details, and upsample the final RGB image without increasing the rendering resolution/cost. Note that from the volume rendering process, we can also get the expected depth for each pixel [29].